OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

do you know what substitution is,

OpenStudy (anonymous):

11

jimthompson5910 (jim_thompson5910):

g(x) = x^2 + 2 g(3) = (3)^2 + 2 ... Replace every 'x' with 3 g(3) = ??? .... Evaluate

OpenStudy (anonymous):

@miquakk please dont give answers